Two angles are complementary. The sum of the measure of the first angle and half the second angle is 72.5. Find the measure of the angles.
what is the measure of the smaller angle?
what is the measure of the other angle?

Answer by stanbon(75887)

Two angles are complementary. The sum of the measure of the first angle and half the second angle is 72.5. Find the measure of the angles.
what is the measure of the smaller angle?
what is the measure of the other angle?
--------------------------------------------
Let one of the angles be "x"
The other angle is "180-x"
---
Equation:
180-x + (x/2) = 72.5
Muliply thru by 2 to get:
2(180-x) + x = 145
360 - 2x + x = 145
-x = - 115
x = 115 (larger angle)
180-x = 65 (smaller angle)
---------------------------------------
Cheers,
Stan H.